OpenStudy (anonymous):

A banana = 1; 6 is represented by an orange and 2 bananas; An orange is worth half a mango. What is the value of two mangos, an orange and a banana? a) 21 b) 24 c) 27 d) 30 e) 33 f) 36

OpenStudy (zyi6):

Do you understand what to do here? at all?

OpenStudy (zyi6):

Here is what you do: banana = 1 6 = 2 banana's and an orange. if you do the math, then 1 + 1 = 2, which is the bananas. if 6 is 2 bananas, and 1 orange, than the orange has to be worth 4. If an orange is half a mango, then the mango is 8. so if we have 2 mangos, an orange, and a banana, this is what you do. 8 + 8 = 16 (math for the mangos) 16 + 4 = 20 (mango + orange) 20 + 1 = 21 (mango and orange + banana.) so the answer is A, 21. Do you understnad now?